OpenStudy (anonymous):

Simplify secθ + secθtan^2θ. A.) 1 B.) cosθ C.) secθ D.) sec^3θ

Nnesha (nnesha):

factor!

OpenStudy (anonymous):

example please.

Nnesha (nnesha):

sec theta = x tan theta = y so you can write it as \[\huge\rm (x+xy^2)\] so what is common in these both terms ?

OpenStudy (anonymous):

x

Nnesha (nnesha):

yes right so take it out\[\huge\rm x(1+y^2)\] now you can replace x and y by sec an tan \[\huge\rm sec \theta (1+ \tan^2)\] 1+tan^2 = what ? ^^^^identity

OpenStudy (anonymous):

i still dont understand whatsoever.

Nnesha (nnesha):

it's same like simple algebra first we have to find GCF(greatest common factor) \[\huge\rm sec \theta + \sec \theta \tan^2 \] sec is common so take it out and divide both terms by common factor write your answer in the parentheses \[\sec \theta (1+ tan^2 \theta )\]